OpenStudy (anonymous):

46 krissywatts Passing through (-1,1) and parallel to the line whose equation is 3x-2y=7. write an equation for the line in point-slope form.

OpenStudy (egenriether):

first solve your equation of the parallel line to get the slope of the line you want (they will be the same since the lines are to be made parallel) . y= -3/2 +7. Now you have the slope (-3/2) Using the two given points in your problem as the points in the point slope form gives (y-1) = (-3/2)*(x - (-1)). Simplyfing gives y = -3/2x - 1/2